# Helix: Distributed Serving of Large Language Models via Max-Flow on Heterogeneous GPUs

### v1

@@ 3.4 Analyzing and Speeding up MILP
As Table 2 and 3 show, the number of variables and constraints in the MILP problem scales linearly with the number of compute nodes and network connections. For large clusters with more than 40 nodes, it may take hours before the MILP solver gives a reasonably good solution. To expedite the MILP solving process for large clusters, we introduce three optimizations. First, we prune some of the slow network connections in the cluster. Evaluation in Sec. 5.8 shows that this effectively reduces the problem size without sacrificing much performance. Second, we hint the MILP solver with solutions found by heuristic methods. Since the problem has an exponential solution space, the MILP solver can only cover a small portion within a limited solving time budget. Using solutions from heuristic methods as starting points to the MILP problem expedites the optimization process, especially for large clusters. Sec. 5.8 shows the necessity of starting from heuristic solutions for large clusters. Finally, we notice that the max serving throughput of a cluster is always bounded by the sum of compute throughput of all compute nodes averaged by the total number of layers. The MILP solver uses this as an early-stopping criterion and stops when it finds a solution that is very close to this upper bound.
@@
A common approach for speeding up MILP problems is to relax them to a linear program (LP) by relaxing the integer variables to be linear variables and obtaining a valid solution to the original problem via methods such as rounding the resulting linear variables. We remark that this approach is not viable for the MILP problem formulation above for model placement. This is because the resulting solution from the LP cannot be easily converted to a valid solution of the original problem. The variables for model placement ($s_{i}$ and $b_{i}^{j}$) decide the edge validity variables $d_{i,j}$, which in turn decides the flow variables $f_{i,j}$. Rounding the non-integral values of model placement variables in the relaxed solution may invalidate some or all network connections and thus drastically changing the max flow of the cluster.
4 Helix Runtime
@@
This section discusses the runtime scheduling of requests in Helix. When the coordinator node receives a new request, it runs Helix’s request scheduler to assign the request a per-request pipeline, which we will introduce in Sec. 4.1. Then the coordinator node sends the request to the first compute node in the pipeline. When a compute node receives a requests, it performs inference on the request using the layers it is responsible for in that pipeline and send the request to the following compute node. Fig. 3 shows an overview of Helix.
4.1 Scheduler Design: Per-Request Pipelines
@@
To infer a request in the cluster, the scheduler needs to assign a pipeline for the request. The pipeline contains a sequence of stages, where each stage specifies a compute node and the layers to infer on the compute node. A valid pipeline must infer each layer of the model in correct order when running the stages sequentially. In Helix, instead of having a group of fixed pipelines and assigning requests to them as previous works [4, 15], we propose a per-request pipeline assignment approach, wherein each request will have its own pipeline. The total number of possible pipelines equals to the number of possible paths from source to sink in the graph abstraction of the cluster. The abundant number of pipelines allows the scheduling to fit the capacity of the compute nodes and network connections better. Our Max-Flow formulation enables us to create the per-request pipelines.
@@
The scheduler in Helix uses the graph abstraction of the cluster with model placement solution from MILP, and its max flow solution to guide request scheduling. The scheduler runs on the coordinator node and uses interleaved weighted round-robin (IWRR) [37] to perform request scheduling. IWRR scheduling takes a list of candidates and their weights as input. When queried for scheduling, it selects a candidate with frequency proportional to its weight. Using IWRR allows us to schedule requests following the max flow of the cluster without creating bursts. Below we present our max-flow-based IWRR scheduling process in detail.
Each node in the cluster (including the coordinator node) is represented by an IWRR instance in Helix’s scheduler. Each IWRR instance’s candidates contain all the nodes connected by valid network connections from the corresponding node. The weight of each candidate equals the flow over the corresponding network connection in the max flow solution obtained in Sec. 3.3. When a new request arrives, the scheduler starts scheduling from the IWRR instance representing the coordinator node. It queries the IWRR instance and gets a compute node $c_{1}$, which will host the first pipeline stage. Since the request has not inferred any layers, the first pipeline stage will infer all layers $c_{1}$ holds. For the second pipeline stage, the scheduler will use the IWRR instance representing $c_{1}$ to determine the compute node $c_{2}$ that hosts the second pipeline stage. Because of partial inference mentioned in Sec. 3.3, the layers held by $c_{1}$ and $c_{2}$ might partially overlap. Therefore, the scheduler will check the overlap and only infer the request on layers previously not inferred in the second stage. The scheduler repeats this process until we form a valid pipeline. Based on the way we construct the graph abstraction of a cluster in Sec. 3, the scheduler can always build a valid pipeline for the cluster. The example in Fig. 4 shows the scheduling of two requests using Helix’s request scheduler.
Figure 4: Numbers on each edge represent the flow over the edge in the max flow of the cluster’s graph abstraction. Helix’s request scheduler uses these numbers as weights to perform interleaved weighted round-robin scheduling to form per-request pipelines. The pipelines used to schedule requests 1 and 2 are shown on the right.
4.2 KV-Cache Estimation
@@
When serving LLMs, each GPU has a limited amount of VRAM to store the KV-cache of requests during inference. If the requests running concurrently on the GPU requires more KV-cache than this limit, the execution engine has to offload some requests to main memory, which significantly harms throughput. Therefore, in the scheduler we maintain an estimation of KV-cache usage of all compute nodes, and mask out compute nodes that exceed high-water mark when running interleaved weighted round-robin. We can schedule more requests to those compute nodes only after some requests currently running on those nodes have finished. This mask ensures that we do not over-subscribe the GPUs’ KV-cache.
5 Evaluation
@@
  • Can Helix provide higher throughput compared to existing systems in different settings (Sec. 5.3)?
  • Does Helix sacrifice latency to obtain the high throughput (Sec. 5.3)?
  • How does network heterogeneity affect Helix’s performance (Sec. 5.4)?
  • How does the degree of heterogeneity affect Helix’s performance (Sec. 5.5)?
  • Why does Helix achieve better performance compared to existing systems (Sec. 5.6 and 5.7)?
5.1 Implementation
Helix prototype
We implemented a multi-replica pipeline parallel system with 1.5k LoC in Python and 1.7k LoC in C++. For model execution engine, we adopt the latest release of vLLM [16] (0.4.0post1) to avoid reimplementing basic LLM inference optimizations like iteration level scheduling and PageAttention. We implemented a pool of Pages unified for all local layers in a node, since requests may only execute a subset of all local layers. For inter-node communication, we use ZeroMQ [38] to transfer both metadata and intermediate tensors. We use Gurobi [10] as the solver of our MILP.
Simulator
Besides the prototype system, we build a event-based simulator for distributed LLM inference in heterogeneous clusters with 14k LoC in Python. We tune the simulator with profiled data collected from the prototype system for a high fidelity. Running simulation is much faster and requires less computation than the real system. It also allows us to explore more diverse settings of network conditions, machine heterogeneity and cluster scale. We test the fidelity of the simulator against the prototype system in Sec. 5.3.
5.2 Experiment Setup
@@
We evaluate Helix on LLaMA [40, 41], a representative and popular open-source Transformer model family. Specifically, we use LLaMA 30B and LLaMA 70B to study the system performance on models of different sizes. We run model inference with half-precision (FP16).
@@
We evaluate with three cluster setups: (1) single cluster, (2) distributed cluster and (3) high GPU heterogeneity. For the single cluster setup, we create a heterogeneous cluster with 4 A100 nodes, 8 L4 nodes and 12 T4 nodes within one region on the Google cloud. Our evaluation focuses on single-GPU nodes, as multi-GPU nodes are much harder to allocate on the cloud [39], while our implementation also works for multi-GPU nodes by leveraging tensor model parallelism across GPUs on the same node as supported by vLLM [16]. Average transmission bandwidth between machines is 10 Gb/s and average latency is smaller than 1ms. The distributed cluster setup has three clusters that contain 4 A100 nodes, 2 L4 nodes + 8 T4 nodes, and 6 L4 nodes + 4 T4 nodes respectively. Intra-cluster network condition mirrors that of single cluster setup. Inter-cluster communication has an average bandwidth of 100 Mb/s and an average latency of 50 ms (from our profiling results on Google Cloud). For high GPU heterogeneity cluster used in simulation, the cluster contains 4 A100 nodes, 6 V100 nodes, 8 L4 nodes, 10 T4 nodes, 4 2$\times$×L4 nodes, 6 2$\times$×T4 nodes and 4 4$\times$×T4 nodes. Multi-GPU nodes run tensor parallelism for the layers assigned to them.
@@
The traces we use come from Azure Conversation dataset [27]. Fig. 5 shows the length distribution and arrival rate of this dataset. We remove requests with input lengths larger than 2048 or output lengths larger than 1024 to fit in the GPUs we use. The pruned dataset contains 16657 requests with an average input length of 763 and an average output length of 232. We have two settings for arrival rates. For online setting, we use real arrival rates from Azure Conversation dataset. We scale the average arrival rate to 75% of the cluster’s peak throughput to avoid bursts of requests leading to OOM in the system. For offline setting, we allow requests to arrive at the rate needed to fully utilize the cluster. This mimics running offline inference on a dataset. We refer to the two settings as online and offline serving.
(a) Input and output length.
(b) Arrival rate.
Figure 5: Length and arrival rates of requests in Azure Conversation dataset.
@@
We allow the MILP solver to search w/ and w/o partial inference and w/ and w/o cluster pruning. We also hint the MILP solver with solutions from Petals / Swarm / separate pipelines. Solving times out when the MILP solver does not find better solutions in 10 minutes. The total search budget for each cluster setup is 4 hours on a 14 core CPU.
@@
To the best of our knowledge, there are no heterogeneous LLM serving systems with model placement and scheduling applicable to our settings. Therefore, we adopt ideas from a heterogeneous LLM training system, Swarm [31], to build a competitive heterogeneous baseline. We also build another baseline that builds on homogeneous pipelines, which run one separate pipeline for each type of machine. We refer to the two baselines as Swarm and separate pipelines (SP). For Swarm, we implement the method in our system because the original system is designed for training and cannot be directly used for inference. Our implementation ensures that requests in the prompt and decode phase always follow the same path, which is crucial for the correctness of inference. We set the number of pipeline stages to the minimum that allows the weakest GPU in the cluster to hold one stage of layers with half its VRAM. This minimizes the pipeline depth and leaves enough VRAM for KV-cache on GPUs, both of which are crucial to performance. For separate pipelines, each pipeline serves one replica of the model and layers are equally distributed among machines within the pipeline.
@@
(a) LLaMA-30B - Online
(b) LLaMA-70B - Online
(c) LLaMA-30B - Offline
(d) LLaMA-70B - Offline
Figure 6: Decode throughput comparison between different methods for online and offline serving in single cluster setup.
(a) LLaMA-30B prompt latency
(b) LLaMA-30B decode latency
(c) LLaMA-70B prompt latency
(d) LLaMA-70B decode latency
Figure 7: Prompt and decode latency comparison of online serving in a single cluster setup. In the figure, the box shows the 25 and 75 percentile, the whisker shows the 10 and 90 percentile, the orange line shows the median, and the green triangle shows the mean.
5.3 Single Cluster
Real system evaluation.
This section evaluates Helix’s prototype system in single cluster setup. We evaluate both LLaMA 30B and LLaMA 70B for online and offline serving, forming four groups of experiments. Each method uses the same model placement plan for online and offline serving. Comparison of decode throughput between different methods for offline and online serving are shown in Fig. 6. Comparison of prompt and decode latency between different methods for online serving is shown in Fig. 7.
Offline serving.
For LLaMA 30B, each type of compute node is able to serve a separate pipeline by themselves. In this case, the best model placement found by Helix’s model placement planner contains three separate pipelines. The overall performance of Helix is very close to separate pipelines, despite the small performance gain from better utilization of KV cache enabled by the KV cache estimation in Sec. 4.2. Swarm’s model placement plan introduces a bottleneck and under-utilizes the A100 nodes, which we will discuss in more detail in Sec. 5.6. As a result, Helix achieves 2.14$\times$ boost in decode throughput compared with Swarm.
For LLaMA(70B), things are much different. When serving one pipeline with each type of machine, the majority of GPU VRAM will be used for model parameters not leaving sufficient space for the KV cache, which greatly limits the decode throughput. Thus, in this case, the decode throughput of separate pipelines degrades significantly. Similar to LLaMA 30B case, Swarm introduces a bottleneck in its model placement plan, causing under-utilization of both A100 and L4 nodes. As a result, Helix achieves 1.94×1.94$\times$1.94 × and 1.86×1.86$\times$1.86 × decode throughput compared with Swarm and separate pipelines baseline.
Online serving.
For LLaMA 30B, Helix has similar prompt and decode latency compared with separate pipelines baseline, since the model placements are identical. On the other hand, Swarm has 47% higher prompt latency and 13% higher decode latency compared with Helix, even though its throughput is just half of Helix (Fig. 6(a)). The reason is that Swarm under-utilizes the A100 nodes, causing most compute to happen on the slow L4 and T4 GPUs, which will be discussed in detail with an example in Sec. 5.6. For LLaMA 70B, running separate pipelines yields the lowest prompt and decode latency. The reason is that the 8 L4 and 12 T4 nodes are heavily bounded by KV cache capacity (too much VRAM used for parameters) and serves at a very low throughput. The majority of requests are served by the 4 A100 nodes, which has very low latency. This however results in severe waste of compute capacity of the L4 and T4 nodes and thus resulting lower throughput as seen in Fig. 6(b).
Simulator fidelity evaluation.
We also perform the same experiments in the single-cluster setup on the simulator mentioned in Sec. 5.1. Then we compare its results with that of real machine experiments. Fig. 6(a) and 6(b) shows that for online serving experiments, in which the cluster runs at 75% peak throughput, the average decode throughput difference between simulation and real evaluation result is less than 1%. Fig. 7 shows the prompt and decode latency comparison for online serving experiments. We find that for all methods, there is a constant difference of $\sim$150ms for both LLaMA 30B and 70B. After profiling, we identify this difference as the CPU-GPU data transfer overhead not modeled in the simulator. Since the difference is consistent for all methods, it does not affect the comparison between them. For offline serving experiments, Fig. 6(c) and 6(d) shows that the simulator reports a consistently higher throughput for all methods. This gap is also caused by the constant CPU-GPU transfer overhead not modeled. Since the difference is consistent between different methods, the simulator still is valuable as it serves the purpose of comparing the performance of different methods.
(a) Online
(b) Offline
Figure 8: Decode throughput comparison for online and offline serving in a distributed cluster setup.
(a) LLaMA-30B
(b) LLaMA-30B
@@ 5.3 Single Cluster
(d) LLaMA-70B
(e) Throughput with more heterogeneity
Figure 9: (a-d) Prompt and decode latency comparison for online serving in the distributed cluster setup. (e) Decode throughput comparison for serving LLaMA 70B in a cluster with 42 nodes and 7 node types.
Figure 10: Left: decode throughput of Helix, Petals and Swarm in single and distributed cluster setup. Right: A case study of the model placement found by the three methods for serving LLaMA 70B in single cluster setup. The number in each node represents the number of layers the node holds.
5.4 Distributed Clusters
In this section, we evaluate Helix with a distributed cluster setup in simulation. Similar to single cluster experiments, we evaluate on both LLaMA 30B and 70B for online and offline serving. Fig. 8 shows the decode throughput comparison. For LLaMA 30B, the decode throughput of all methods drop by 10-20% compared with single cluster settings. For LLaMA 70B, decode throughput of all methods drop by half. This comes from the slow inter-cluster network connections, and larger models are more affected by this because they require a deeper serving pipeline. For LLaMA 70B, which is heavily affected by the slow network, we find that the max pipeline depth of Helix’s model placement is 28% shorter than that of Swarm’s, because Helix’s MILP model placement planners finds better placement strategies to load more layers into L4 and A100 machines to reduce network communication in this case. By comparing the model placement of Helix for LLaMA 70B in single and distributed cluster setups (which only differs in network condition), the model placement for distributed cluster reduces pipeline depth by 19%. This demonstrates that Helix’s model placement takes network condition into account, which makes a 2.0$\times$ and 1.5$\times$ improvement over Swarm and separate pipelines.
Fig. 9 shows the prompt and decode latency comparison for online serving. For LLaMA 30b, Helix has a similar latency as separate pipelines since both methods use the same model placement in this case. Swarm, on the other hand, has a much higher latency for the same reason as Sec. 5.3. For LLaMA 70B, separate pipelines has the lowest latency, because the 12 slow T4 nodes are bounded by memory and contribute to only a quarter of the total throughput. The under-utilization of T4’s compute power improves decode latency by 28% at the cost of 45% decode throughput. For Swarm, we observe severe congestion when the cluster is running at peak throughput. The average prompt latency reaches 70s, which is 7.5$\times$ that of Helix’s. We will discuss more about the congestion of Swarm with a case study in Sec. 5.7.
5.5 Cluster Heterogeneity

### v2

@@ 4.5. Analyzing and Speeding up MILP
As Table 5 and 6 show, the number of variables and constraints in the MILP problem scales linearly with the number of compute nodes and network connections. For large clusters with more than 40 nodes, it may still take hours before the MILP solver gives a reasonably good solution. To expedite the MILP solving process for large clusters, we introduce three optimizations. First, we prune some of the slow network connections in the cluster. Evaluation in Sec. 6.8 shows that this effectively reduces the problem size without sacrificing much performance. Second, we hint the MILP solver with solutions found by heuristic methods. Since the problem has an exponential solution space, the MILP solver can only cover a small portion within a limited solving time budget. Using solutions from heuristic methods as starting points for the MILP problem expedites the optimization process, especially for large clusters. Sec. 6.8 shows the necessity of starting from heuristic solutions for large clusters. Finally, we notice that the max serving throughput of a cluster is always bounded by the sum of compute throughput of all compute nodes averaged by the total number of layers. The MILP solver uses this as an early stop criterion and stops when it finds a solution that is very close to this upper bound. We remark that, for further scaling of Helix to hundreds or even thousands of nodes, one viable approach is to first partition the nodes into multiple smaller clusters using heuristics and then apply Helix independently.
4.6. Replacing MILP with LP or Heuristics?
@@
A common approach for speeding up MILP problems is to relax them to a linear program (LP) by relaxing the integer variables to be linear variables and obtaining a valid solution to the original problem via methods such as rounding the resulting linear variables. We remark that this approach is not viable for the MILP problem above. This is because the resulting solution from the LP cannot be easily converted to a valid solution of the original problem. The variables for model placement ($s_{i}$ and $b_{i}^{j}$) decide the edge validity variables $d_{i,j}$, which in turn decides the flow variables $f_{i,j}$. Rounding the non-integral values of model placement variables in the relaxed solution may invalidate some or all network connections and thus drastically changing the max flow.
Our preliminary exploration also indicates that using simpler heuristics cannot guarantee good performance across various cluster setups, because of the exponential solution space of this problem. Sec. 6.6 shows that the model placement found by Helix with MILP is much better than that of the heuristic baselines.
Figure 4. Topology graph of a cluster, where each vertex is a compute node, and each edge is a valid network connection. Numbers over edges represent the flow over the network connection in the max flow solution. The pipelines used to schedule requests 1 and 2 are shown on the right.
[…]
Figure for IWRR.
5. Helix Runtime
@@
This section discusses the runtime scheduling of requests in Helix. When the coordinator node receives a new request, it runs Helix’s request scheduler to assign the request a per-request pipeline, which we will introduce in Sec. 5.1. Then the coordinator node sends the request to the first compute node in the pipeline. When a compute node receives a requests, it performs inference on the request using the layers it is responsible for in that pipeline and sends the request to the following compute node. Fig. 3 shows the overview of Helix.
5.1. Scheduler Design: Per-Request Pipelines
@@
To infer a request in the cluster, the scheduler needs to assign a pipeline for the request. The pipeline contains a sequence of stages, where each stage specifies a compute node and the layers to infer on the compute node. A valid pipeline must infer each layer of the model exactly-once and in correct order when running the stages sequentially. Existing works (borzunov2022petals, ; jiang2023hexgen, ) use fixed pipelines, in which each pipeline contains a disjoint set of machines, and assign requests to those pipelines. In Helix, instead of using fixed-pipelines, we propose a per-request pipeline assignment approach, wherein each request will have its own pipeline and the pipelines may intersect with each other. The total number of possible pipelines equals the number of possible paths from source to sink in the graph abstraction of the cluster. The abundant number of pipelines allows the scheduling to better fit the capacity of the compute nodes and network connections. Our Max-Flow formulation enables us to create the per-request pipelines.
@@
The Helix request scheduler performs scheduling based on the cluster’s topology graph (Fig. 4). In the topology graph, vertices correspond to the nodes in the cluster. Directed edges correspond to the valid network connections (under the model placement found by solving the MILP in Sec. 4). We bind an interleaved weighted round-robin (IWRR) (tabatabaee2021interleaved, ) scheduler to each vertex. The IWRR scheduler takes as input a list of candidates and their weights. To schedule a request, it selects a candidate with frequency proportional to its weight. For each vertex $u$, the IWRR scheduler’s candidates contain all vertices $v$ such that directed edge $(u,v)$ exists (i.e. $(u,v)$ represents a valid network connection). The weight of $v$ equals the flow over the network connection of $(u,v)$ in the max flow solution. Using IWRR allows us to schedule requests following the max flow without creating bursts.
Helix’s request scheduler runs on the coordinator node. When a new request arrives, it first uses the IWRR scheduler of the vertex representing the coordinator node to determine the compute node $c_{1}$ for the first pipeline stage. Then, it uses the IWRR scheduler of the vertex representing $c_{1}$ to determine the compute node $c_{2}$ for the second stage, and repeats this process until a valid pipeline is established. Fig. 4 shows the scheduling of two requests. After setting the request’s pipeline, the coordinator node sends the request to the first compute node in the pipeline to begin inference. During inference, Helix adopts a dynamic batching strategy where each node includes all requests received during the processing of the previous batch to form a new one. This best-effort batching occurs without additional waiting periods.
5.2. KV-Cache Estimation
@@
When serving LLMs, each GPU has a limited amount of VRAM to store the KV-cache of requests during inference. If the requests running concurrently on the GPU require more KV-cache than this limit, the execution engine has to offload some requests to main memory, which significantly harms throughput. However, we do not know exactly how much KV-cache each request will use because the length of the output is unknown before inference finishes. Therefore, in the scheduler we maintain an estimation of KV-cache usage of all compute nodes using average output length, and mask out compute nodes that exceed the high water mark when running IWRR. We can schedule more requests to the compute nodes only after some requests currently running on those nodes have finished. This mask ensures that we do not oversubscribe the GPU’s KV-cache.
(a) Length distribution.
(b) Arrival rate.
Figure 5. Statistics of Azure Conversation dataset.
[…]
Figure for Azure Conversion dataset.
6. Evaluation
Table 7. Network bandwidth between machines in asia-east, us-central, europe-west and australia-southeast. Measured with iperf3 on Google Compute Engine.
Receiver Sender
asia-east2-a us-central1-f eu-west3-c au-se1-c
asia-east2-a / 123 Mbps 67 Mbps 175 Mbps
us-central1-f 122 Mbps / 204 Mbps 123 Mbps
eu-west3-c 61 Mbps 196 Mbps / 54 Mbps
au-se1-c 159 Mbps 118 Mbps 63 Mbps /
(a) LLaMA-30B - Offline
(b) LLaMA-30B - Online
(c) LLaMA-70B - Offline
(d) LLaMA-70B - Online
(e) LLaMA-30B - Prompt
(f) LLaMA-30B - Decode
(g) LLaMA-70B - Prompt
(h) LLaMA-70B - Decode
Figure 6. Single cluster results: (a - d) decode throughput, (e - h) prompt and decode latency for online serving. The box shows the 25 and 75 percentile. The whisker shows the 5 and 95 percentile. The red line shows median.
[…]
Figure for single cluster results.
@@ 6. Evaluation
  • Can Helix provide higher throughput for heterogeneous GPUs in a single cluster? (Sec. 6.3)
  • Does Helix sacrifice latency for throughput? (Sec. 6.3)
  • Can Helix provide higher throughput for heterogeneous GPUs in geo-distributed clusters? (Sec. 6.4)
  • Can Helix provide consistently high throughput when the degree of GPU heterogeneity increases? (Sec. 6.5)
  • Why does Helix achieve better performance compared to existing systems (Sec. 6.6 and 6.7)?
6.1. Implementation
Helix prototype.
We implemented a multi-replica pipeline parallel system with 1.5k LoC in Python and 1.7k LoC in C++. For model execution, we adopt the latest release of vLLM (kwon2023vllm, ) (0.4.0post1) to avoid re-implementing basic LLM inference optimizations. We implemented a unified page pool atop vLLM to support partial inference. We use ZeroMQ (zmq, ) for inter-node communication and Gurobi (gurobi, ) as MILP solver.
Simulator.
We also implemented a simulator for distributed LLM inference with 14k LoC in Python. It supports simulation for heterogeneous GPUs and network conditions. It gives us the flexibility to explore more diverse settings for network, GPU heterogeneity and cluster scale. Sec. 6.3 evaluates the fidelity of the simulator and shows that the simulation errors are less than 5% for all metrics.
(a) LLaMA-30B
(b) LLaMA-70B
(c) LLaMA-30B
(d) LLaMA-30B
(e) LLaMA-70B
(f) LLaMA-70B
Figure 7. Geo-distributed clusters: (a - b) decode throughput, (c - f) prompt and decode latency for online serving. ”H” stands for Helix, ”S” stands for Swarm. Box: 25 and 75 percentile. Whisker: 5 and 95 percentile. Red line: median.
[…]
Figure for geo-distributed cluster results.
6.2. Experiment Setup
@@
We evaluate Helix on LLaMA (touvron2023llama, ; touvron2023llama2, ), a representative and popular open-source Transformer model family. Specifically, we use LLaMA-1 30B and LLaMA-2 70B to study the system performance on models of different sizes. We run model inference with half-precision (FP16). In subsequent sections, we refer to these two models as LLaMA 30B and LLaMA 70B.
@@
We evaluate with three cluster setups: (1) single cluster (Sec. 6.3), (2) geo-distributed clusters (Sec. 6.4), and (3) high GPU heterogeneity cluster (Sec. 6.5). First, the single cluster setup contains 4 A100 nodes, 8 L4 nodes and 12 T4 nodes connected by 10Gb/s network. We configure the network bandwidth to 10 Gb/s, as this rate is sufficient to ensure that LLM serving is limited by GPU throughput rather than network capacity. We allocate the nodes within one region on the Google Cloud. Second, the geo-distributed clusters contain three clusters with (i) 4 A100 nodes, (ii) 2 L4 nodes + 8 T4 nodes, and (iii) 6 L4 nodes + 4 T4 nodes. Inter-cluster communication has an average bandwidth of 100 Mb/s and an average latency of 50 ms. We configure the bandwidth to 100 Mb/s to simulate cross-region network limitations, based on our profiling results in Table 7. Finally, the high GPU heterogeneity cluster contains 4 A100 nodes, 6 V100 nodes, 8 L4 nodes, 10 T4 nodes, 4 2$\times$×L4 nodes, 6 2$\times$×T4 nodes and 4 4$\times$×T4 nodes. The latter two setups are performed in simulation. Our evaluation of the prototype system focuses on single-GPU nodes, as multi-GPU nodes are much more difficult to allocate in the cloud (thorpe2023bamboo, ). Our implementation also works for multi-GPU nodes by leveraging tensor model parallelism across GPUs on the same node as supported by vLLM (kwon2023vllm, ).
@@
The traces we use come from Azure Conversation dataset (patel2023splitwise, ). Fig. 5 shows the length distribution and arrival rate of this dataset. We remove requests with input lengths larger than 2048 or output lengths larger than 1024 to maintain reasonable runtime memory usage for vLLM. The pruned dataset contains 16657 requests with an average input length of 763 and an average output length of 232. We remark that this dataset is more challenging than the ones used by prior work (kwon2023vllm, ) due to longer request length. We have two settings for arrival rates. For online setting, we use real arrival rates from Azure Conversation dataset. We scale the average arrival rate to 75% of the cluster’s peak throughput to avoid bursts of requests leading to OOM in the system. For offline setting, we allow requests to arrive at the rate needed to fully utilize the cluster. This mimics running offline inference on a dataset. We refer to the two settings as online and offline serving.
@@
We allow the MILP solver to search w/ and w/o partial inference and cluster pruning. We also hint the MILP solver with solutions from Petals / Swarm / separate pipelines. Solving times out when the MILP solver does not find better solutions in 10 minutes. The max search budget for each cluster setup is 4 hours on a 14-core CPU.
@@
To the best of our knowledge, there are no heterogeneous LLM serving systems with model placement and scheduling applicable to our settings. Therefore, we adopt ideas from a heterogeneous LLM training system, Swarm (ryabinin2023swarm, ), to build a competitive heterogeneous baseline. We also build another baseline that handles GPU heterogeneity by serving one model replica with each type of machine. We refer to the two baselines as Swarm and separate pipelines (SP). To ensure a fair comparison, we implemented the SP baseline within our system rather than utilizing other systems. We further compare with the model placement of Petals (borzunov2022petals, ) in Sec. 6.6, which is a decentralized LLM serving system for volunteer computing. We do not compare with Petals in end-to-end serving as it lacks centralized request scheduling. For Swarm, we implemented their model placement and scheduling algorithm in our system, since their original system can not be used for inference. We set the number of pipeline stages to the minimum that allows the weakest GPU to hold one stage with half its VRAM. This minimizes the pipeline depth and leaves enough VRAM for KV-cache, both of which are crucial to performance.
@@
(a) LLaMA-70B
(b) LLaMA-70B
@@
Figure 8. High GPU-heterogeneity clusters setup: (a) decode throughput, (b - c) prompt and decode latency for online serving. ”H” stands for Helix, ”S” stands for Swarm.
[…]
Figure for high GPU-heterogeneity cluster results.
(a) Decode throughput.
(b) Model placement case study. Numbers on each node represent the number of layers the node holds.
Figure 9. (a) Comparing different model placement methods with offline serving of LLaMA 70B. (b) The case study shows model placements for serving LLaMA 70B on the single cluster setup.
[…]
Figure for model placement.
6.3. Single Cluster
This section evaluates Helix with online and offline serving of LLaMA 30B and 70B in the single cluster setup. Fig. 6 shows the results. For LLaMA 30B, each GPU type has enough nodes to serve at least one individual pipeline. The best model placement discovered by Helix serves three replicas of the model, each with one type of GPU. As a result, Helix and SP achieve similar performance: Helix has 4% and 14% higher decode throughput for offline and online serving because of its better KV-cache utilization; the prompt latency is 2% lower and decode latency is 10% higher. Compared with Swarm, Helix achieves 2.14$\times$ and 2.07$\times$ higher decode throughput for offline and online serving. Helix also reduces prompt and decode latency by 32% and 12% for online serving. We find that Swarm’s model placement introduces a bottleneck and under-utilizes the A100 nodes, which we discuss in more detail in Sec. 6.6. We note that the comparatively higher latency observed in our experiments, relative to other LLM serving systems (kwon2023vllm, ; patel2023splitwise, ), can be attributed to our use of less powerful GPUs (T4 and L4) in contrast to the A100 and H100 GPUs utilized in other studies.
For LLaMA 70B, nodes from a single GPU type can not serve one model replica by themselves while leaving enough VRAM for KV-cache. In this case, SP’s throughput significantly decreases. Compared with SP, Helix achieves 1.86$\times$ and 1.69$\times$ higher decode throughput in offline and online serving. The average latency of Helix is higher than SP because SP serves majority of requests with A100 nodes and under-utilizes the L4 and T4 nodes. Compared with Swarm, Helix achieves 1.94$\times$ and 2.00$\times$ higher decode throughput in offline and online serving. Helix’s prompt latency is 15% lower, but decode latency is 16% higher because of Helix’s high GPU utilization. Similar to LLaMA 30B, Swarm under-utilizes A100 nodes because of its model placement.
We also conduct these experiments in the simulator (see Sec. 6.1) to examine its fidelity. The simulation results are shown alongside the real system evaluation results in Fig. 6. The average error of decode throughput is lower than 5%. The average error of prompt and decode latency is lower than 5% and 4% respectively. This indicates that our simulator achieves high fidelity and serves the purpose of comparing the performance of different methods.
6.4. Geo-Distributed Clusters
This section evaluates Helix for both online and offline serving of LLaMA 30B and 70B in geo-distributed clusters using our high-fidelity simulator in Sec. 6.1. Fig. 7 shows the results. Although this setup involves the same set of GPUs as the single cluster, the slow inter-cluster network causes all methods to have lower throughput and higher latency.
For LLaMA 30B, the best model placement found by Helix still consists of three pipelines served by three types of machines separately. Compared with SP, Helix has 7% and 10% higher decode throughput for offline and online serving. The prompt latency is 14% lower and decode latency is 2% higher. Compared with Swarm, Helix achieves 2.41$\times$ and 2.33$\times$ higher decode throughput for both online and offline serving. Helix also reduces prompt and decode latency by 66% and 24% for online serving.
On the other hand, serving LLaMA 70B is more sensitive to slow network because of larger activation size and model depth. Helix reduces network overhead by using model placements with shallower pipeline depth. The model placement found by Helix reduces pipeline depth by 28% compared to Swarm. It is also 19% shallower than the one found by Helix when network is fast. Helix’s model placement planner balances network overhead with single node’s GPU utilization, achieving high throughput and low latency in geo-distributed GPU clusters. Compared with SP, Helix achieves 1.61$\times$ and 1.79$\times$ higher decode throughput for offline and online serving, respectively. SP has lower average latency because it serves most requests with fast GPUs, while 12 T4 nodes are under-utilized because of VRAM constraints. Compared with Swarm, Helix achieves 1.92$\times$ and 1.97$\times$ higher decode throughput for offline and online serving. Helix also reduces prompt and decode latency by 21% and 7%. We observe severe congestion during offline serving with Swarm. The average prompt latency reaches 71s, which is 7.5$\times$ that of Helix’s. We will discuss more about the congestion of Swarm with a case study in Sec. 6.7.
6.5. High GPU-Heterogeneity Cluster